what is the future value in 3.1998 years pf a Present value pf $0.86 continuously compounded at 17.8884%
Answer #1

We use the formula:  
A=P(e)^rn
where
A=future value
P=present value  
r=rate of interest
n=time period.

e=2.71828

A=0.86*(2.71828)^(0.178884*3.1998)

=0.86*1.77250294

=$1.52(Approx)